OpenStudy (anonymous):

Question 1 (Multiple Choice Worth 1 points) (03.05) A science experiment begins with a metal at -100° Celsius. The following function describes the temperature change per minute: f(x) = 89x - 100°. How will the graph of this function change if the metal is at 25° at the start of the experiment? The line will shift vertically up by 125°. The line will shift vertically up by 25°. The line will shift vertically down by 125°. The line will shift vertically down by 25°. Question 2 (Fill-In-The-Blank Worth 1 points) (03.05) Given f(x) and g(x) = f(x) + k, look at the graph bel

OpenStudy (anonymous):

Question 1 (Multiple Choice Worth 1 points) (03.05) A science experiment begins with a metal at -100° Celsius. The following function describes the temperature change per minute: f(x) = 89x - 100°. How will the graph of this function change if the metal is at 25° at the start of the experiment? The line will shift vertically up by 125°. The line will shift vertically up by 25°. The line will shift vertically down by 125°. The line will shift vertically down by 25°. Question 2 (Fill-In-The-Blank Worth 1 points) (03.05) Given f(x) and g(x) = f(x) + k, look at the graph below and determine the value of k. k = ______ Graph of f of x and g of x. f of x equals 1 over 3 x minus 2 and g of x equals 1 over 3 x plus 3. Numerical Answers Expected! Answer for Blank 1: Question 3 (Multiple Choice Worth 1 points) (03.05) A movie rental store charges a membership fee of $10.00 plus $1.50 per movie rented. What is a possible domain that shows the cost to a member of renting movies? The domain is all integers. The domain is all even integers. The domain is all integers 0 or greater. The domain is all integers 10 or greater. Question 4 (Multiple Choice Worth 1 points) (03.05) You attend an amusement park with your family. Your parents buy you an all-ride pass for $20, shown as f(x). Instead of getting a pass, your parents decide to pay $4 for each ride they take, shown as g(x). What function shows the correct combination of these two functions to represent the total cost to them of attending the amusement park that day, shown as h(x)? f(x) = 20x, g(x) = 4, h(x) = 20x + 4 f(x) = 20, g(x) = 4, h(x) = 4 + 20 f(x) = 20, g(x) = 4x, h(x) = 4x + 20 f(x) = 20x, g(x) = 4x, h(x) = 20x + 4x Question 5 (Multiple Choice Worth 2 points) (03.05) A student is training for the track team during the first three months of the year by running for 30 minutes every day, while averaging 5 miles each day. Select the graph that correctly shows the total miles ran during the time of his training program. Graph of miles versus hours. Line segment between points zero comma zero and 450 comma 45. Graph of hours versus miles. Line segment between points zero comma zero and 45 comma 450. Graph of hours versus miles. Ray starting at point zero comma zero and passing through 45 comma 450. Graph of miles versus hours. Ray starting at point zero comma zero and passing through 450 comma 45. Question 6 (Multiple Choice Worth 1 points) (03.05) A cell phone plan has a monthly cost that is shown in the table below. What is the correct statement regarding the average rate of change during the 40-minute time of talk? Total minutes of talk time Monthly cost of cell phone 0 $19.95 10 $20.45 20 $20.95 30 $21.45 40 $21.95 The average rate of change is $0.50, meaning that for each minute of talk time, the monthly bill increases by $0.50. The average rate of change is $0.50, meaning that for every ten minutes of talk time, the bill increases by $0.50. The average rate of change is $0.05, meaning that for each minute of talk time, the monthly bill increases by $0.05. The average rate of change is $0.05, meaning that for every ten minutes of talk time, the bill increases by $0.05. Question 7 (Multiple Choice Worth 1 points) (03.05) The graph shows the temperature of ice cream in an ice chest during an overnight trip. What is the domain of this function? Graph showing a line segment that goes from ordered pairs 0, negative 10 to 30, 30. The domain is negative 10 through 30. The domain is all integers from negative 10 through 30. The domain is all real numbers 0 through 30. The domain is all integers from 0 through 30. Question 8 (Multiple Choice Worth 1 points) (03.05) A student is running a 5-kilometer race. He runs 1 kilometer every 3 minutes. Select the function that describes his distance from the finish line. f(x) = - 1 over 3 x + 5 f(x) = - 1 over 5 x + 3 f(x) = 1 over 5 x + 5 f(x) = 1 over 3 x + 3 Question 9 (Multiple Choice Worth 1 points) (03.05) Your friend has $80 when he goes to the fair. He spends $4 to enter the fair and $12 on food. Rides at the fair cost $1.25 per ride. Which function can be used to determine how much money he has left over after x rides? f(x) = 1.25x + 64 f(x) = -16x + $80 f(x) = -1.25x + 64 f(x) = -1.25x - 64

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

@freckles @Hero

OpenStudy (freckles):

This is a lot to look at at one time.

OpenStudy (freckles):

I will help you with the 1st question.

OpenStudy (anonymous):

sorry do you want me to separate them and open up a new question but can you jut do the first one first and i will tag you agin and separate the questions

OpenStudy (freckles):

if we started at -100 the equation gives f(x)=89x-100 so if we start at 25 wouldn't the equation be f(x)=89x+25

OpenStudy (freckles):

so find the difference of 25 and -100

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-125

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

is 25 greater than -100? and the first number was -100 and the second number was 25 so did we move up or down in numbers?

OpenStudy (anonymous):

up?

OpenStudy (freckles):

yes -100 <25

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@freckels

OpenStudy (anonymous):

so the line will shift vertically up for 125

OpenStudy (freckles):

yes

OpenStudy (anonymous):

so a

OpenStudy (freckles):

 yep